There are 2 vats of milk, one counting 25% protein and the other 40% protein. These two are to be mixed in order to get 12 liters of a 35% protein content milk. How many liters of each must be used?
Answer by Boreal(15235)

x liters of 25%
12-x liters of 40%
.25x+(12-x)(.40)=12*(0.35)
0.25x+4.8-0.40x=4.20
-0.15x= -0.6
x=4 liters of 25%
12-x=8 liters of 40%
To go from 25% to 35% requires twice as much 40% (5 % more) compared to 25% (10% less)
.25*4+.40*8=1+3.20=4.20